# Leibniz rule for wedge product in discrete exterior calculus on general polygonal meshes

### v3

@@ 1 Introduction
In this paper, we prove that the wedge product on general polygonal meshes, first introduced in Ptackova and Velho (2017) and later used in applications in Ptackova and Velho (2021), satisfies the Leibniz product rule with discrete exterior derivative (coboundary operator). Furthermore, we show that it corresponds to a cup product of cochains on orientable 2–dimensional pseudomanifolds. We thus partially extend the findings of Schubel et al. (2024) from simplicial to general polygonal setting.
On smooth manifolds, the wedge product allows for building higher degree forms from lower degree ones. Similarly a cup product is a product of two cochains of arbitrary degree $p$ and $q$ that returns a cochain of degree $p+q$ located on $(p+q)$–dimensional cells. The cup product was introduced by J. W. Alexander, E. Čech, and H. Whitney (Whitney, 1957) in 1930’s and it became a well–studied notion in algebraic topology, mainly in the simplicial setting (Munkres, 1984; Fenn, 1983). Later, the cup product was extended also to $n$–cubes (Massey, 1991). In (Arnold, 2012) various cup products on simplicial and cubical $n$–pseudomanifolds are presented. Arnold (2012) also defines L2 cubical Whitney forms and a cubical cup product that fit together with the wedge product of cubical Whitney forms.
In (Gonzalez-Diaz et al., 2011a) cup products on cubical complexes have been applied to compute cohomology rings of 3D digital images (represented as cubical complexes $Q(I)$), thus simplify the combinatorial structure of $Q(I)$ and obtain a homeomorphic cellular complex $P(I)$ with fewer cells, which improves the computational efficiency. Gonzalez-Diaz et al. (2011a) think of the cup product as the way the $n$–dimensional holes obtained in homology are related to each other. It is known that two objects with non isomorphic cohomology rings are not homotopic.
@@ 3 Discrete setting
    Every $(n-1)$-cell is a face of at most two $n$-cells.
@@
    such that $e^{n}_{a}=e_{0}^{n},\,e^{n}_{b}=e^{n}_{k}$, and $e^{n}_{i-1}$ and $e^{n}_{i}$ have a common $(n-1)$-dimensional face ($i=1,\dots,k$).
If in the condition (2) of above definition, every $(n-1)$-dimensional cell is a face of exactly two $n$-cells, then we talk about pseudomanifold without boundary ((Massey, 1991, Chapter IX)). On the other hand, if there exist at least one $(n-1)$-cell which is incident with only one $n$-cell, we talk about pseudomanifold with boundary.
@@
Let $K=\{K^{n}\}$ be a pseudomanifold on the topological space $X$. And let $[e_{\lambda}^{n}:e_{\mu}^{n-1}]$ be the incidence number of the cells $e_{\lambda}^{n}$ and $e_{\mu}^{n-1}$, for $n>0$, such that:
@@
    If $e_{\mu}^{n-1}$ is not a face of $e_{\lambda}^{n}$, then $[e_{\lambda}^{n}:e_{\mu}^{n-1}]=0$.
@@
    If $e_{\mu}^{n-1}$ is a face of $e_{\lambda}^{n}$, then $[e_{\lambda}^{n}:e_{\mu}^{n-1}]=\pm 1$.
@@
    If $e_{\alpha}^{0}$ and $e_{\beta}^{0}$ are the two vertices of the 1-cell $e_{\lambda}^{1}$, then $[e_{\lambda}^{1}:e_{\alpha}^{0}]+[e_{\lambda}^{1}:e_{\beta}^{0}]=0$.
@@
    Let $e_{\lambda}^{n}$ and $e_{\rho}^{n-2}$ be cells such that $e_{\rho}^{n-2}\prec e_{\lambda}^{n}$; let $e_{\alpha}^{n-1}$ and $e_{\beta}^{n-1}$ denote the unique $(n-1)$-cells $e^{n-1}$ such that $e_{\rho}^{n-2}\prec e^{n-1}\prec e_{\lambda}^{n-1}$. Then
    $[e_{\lambda}^{n}:e_{\alpha}^{n-1}][e_{\alpha}^{n-1}:e_{\rho}^{n-2}]+[e_{\lambda}^{n}:e_{\beta}^{n-1}][e_{\beta}^{n-1}:e_{\rho}^{n-2}]=0.$
@@
Let $K$ be an $n$-dimensional pseudomanifold (possibly with or without boundary), and let $e_{1}^{n},e_{2}^{n}$ be $n$-cells with a common $(n-1)$-cell $e^{n-1}$. We define orientations for $e_{1}^{n}$ and $e_{2}^{n}$ to be coherent (with respect to $e^{n-1}$) if:
$[e_{1}^{n}:e^{n-1}]+[e_{2}^{n}:e^{n-1}]=0.$
A set of orientations for all the $n$-cells of $K$ is said to be coherent if it is coherent in the above sense for any pair of $n$-cells with a common $(n-1)$-face.
$K$ is said to be orientable if all its $n$-cells can be oriented such that any pair of $n$-cells sharing an $(n-1)$-dimensional face are oriented coherently. Otherwise it is called nonorientable.
@@
$0\longrightarrow C_{n}(K)\overset{\partial_{n}}{\longrightarrow}\cdots\overset{\partial_{k+1}}{\longrightarrow}C_{k}(K)\overset{\partial_{k}}{\longrightarrow}\cdots\overset{\partial_{1}}{\longrightarrow}C_{0}(K)\longrightarrow 0$
@@ 4 Discrete Differential Forms and the Exterior Derivative
Figure 1: Boundary homomorphism and exterior derivative on polygonal complex. In this figure, our pseudomanifold $K$ has 2–chain $C_{2}=(f_{0},f_{1})$, 1–chain $C_{1}=(e_{0},\dots,e_{5}$), and 0–chain $C_{0}=(v_{0},\dots,v_{4})$. If 2–form $\gamma^{2}$ was given as the exterior derivative of 1–form $\beta^{1}$, then $\gamma=d\beta$ would be defined per faces $f_{0},f_{1}$ by equation (2). Concretely $\gamma(f_{1})=d\beta(f_{1})$ is:
@@
$0\overset{d_{n}}{\longleftarrow}C^{n}(K)\overset{d_{n-1}}{\longleftarrow}\cdots\overset{d_{k}}{\longleftarrow}C^{k}(K)\overset{d_{k-1}}{\longleftarrow}\cdots\overset{d_{0}}{\longleftarrow}C^{0}(K)\longleftarrow 0$
A $q$–form $\alpha$ is said to be a closed form (cocycle) if $d\alpha=0$, and it is called an exact form (coboundary) if there exist a $(q-1)$–form $\beta$ such that $d\beta=\alpha$.
Note that each exact $q$-form $\alpha=d\beta$ is closed, since $d\alpha=dd\beta=0$. Two closed $q$-forms are cohomologous if they differ by an exact $q$–form. We denote cohomology classes by $H^{\ast}$:
@@ 5 Wedge and Cup Products
A differential $k$-form $\alpha$ on a smooth $n$–manifold $M$ is a tensor field of type $(0,k)$ that is completely antisymmetric, i.e., for $\alpha\in\Omega^{k}(M)\subset T^{0}_{k}(M)$ we have $\alpha(v,w)=-\alpha(w,v)$ for all $v,w\in TM$. Let $\{e_{1},\dots,e_{n}\}$ be a basis of the tangent space $T_{x}M$ to $M$ at point $x\in M$. Let $\alpha\in T^{0}_{k}(M)$ and $\beta\in T^{0}_{l}(M)$, we define their wedge product $\alpha\wedge\beta\in T^{0}_{k+l}(M)$ at the point $x\in M$ by
@@ 5.1 Simplicial and cubical cup products
Let us first recall the definition of simplicial Whitney forms. Let $c=(v_{0},\dots,v_{p})$ be a $p$–simplex with barycentric coordinates $x_{0},\dots,x_{p}$. Thus $0\leq x_{j}\leq 1\;\forall j$ and $\sum_{i=0}^{p}=1$. The simplicial Whitney $p$–form corresponding to the simplex $c$ is
@@
where $c$ is an $(p+q)$–simplex and $W$ is the simplicial Whitney form.
@@
Let $(v_{0},\dots,v_{k})$ be a $k$–simplex of a pseudomanifold $K$, the simplicial cup product $\cup:C^{p}(K,\mathbb{R})\times C^{q}(K,\mathbb{R})\rightarrow C^{p+q}(K,\mathbb{R})$, i.e., the cup product defined over each simplex $(v_{0},\dots,v_{k})$ reads:
@@
Remark.
In our derivation we have actually obtained $(\alpha^{0}\cup\beta^{2})=\frac{1}{6}(\alpha(v_{0})+\alpha(v_{1})+\alpha(v_{2}))\beta(v_{0},v_{1},v_{2})$ and not $\frac{1}{3}(\alpha(v_{0})+\alpha(v_{1})+\alpha(v_{2}))\beta(v_{0},v_{1},v_{2})$. But in the case of $\frac{1}{6}$, the Leibniz rule for $d(\alpha^{0}\cup\beta^{1})$ is only satisfied if $d\beta=0$, i.e., $\beta$ is a closed form. Whereas for $\frac{1}{3}$, the Leibniz rule for $d(\alpha^{0}\cup\beta^{1})=d\alpha\cup\beta+\alpha\cup d\beta$ is satisfied for any $\alpha$ and $\beta$.
@@
In (Arnold, 2012, Definition 3.2.1) the author defines a cubical cup product $\cup_{c}$, for which the product of two cochains agrees with the wedge product of their cubical Whitney forms (see (Arnold, 2012, Section 3.2.2)), thus the Whitney map provides a connection of cubical cohomology with de Rham cohomology ((Arnold, 2012, Theorem 3.2.12)). From (Arnold, 2012, Definition 3.2.1) we computed the cup product on $k$–dimensional cubes for $0\leq k\leq 2$ and we present the formulas in the following definition:
@@
The cubical cup product $\cup:C^{p}(K,\mathbb{R})\times C^{q}(K,\mathbb{R})\rightarrow C^{p+q}(K,\mathbb{R})$ on a quadrilateral pseudomanifold $K$ is defined by its action on a $k$–dimensional cube $(v_{0},\dots,v_{2^{k}-1})$ as:
@@ 5.2 The cup product on polygons
As advertised earlier, we can extend the definitions of the discrete wedge product to any $p$–polygon such that it still satisfies the defining properties of a cup product (Definition 5.1) and therefore automatically also the properties of Proposition 5.2. Moreover, we show that our polygonal cup product is a skew–commutative bilinear operation on any polygonal forms, not just on closed ones, viz. the Proposition 5.3.
Let us remember that a discrete $q$–form $\alpha^{q}$ is a $q$–cochain defined per each element of a $q$–chain $C_{q}(K)$:
@@
Furthermore, we sum two discrete $q$–forms $\alpha^{q},\beta^{q}$ by summing their values per each element of a $q$–chain $C_{q}(K)$ and multiply a $q$–form $\alpha^{q}$ by a scalar $t\in\mathbb{R}$ in element–wise fashion too:
@@
Let $K$ be a 2–pseudomanifold whose 2–cells (faces) are polygons. The cup product $C^{k}(K,\mathbb{R})\times C^{l}(K,\mathbb{R})\rightarrow C^{k+l}(K,\mathbb{R})$ of two discrete forms $\alpha^{k},\beta^{l}$ is a $(k+l)$–form $\alpha^{k}\cup\beta^{l}$ defined on each $(k+l)$–cell of $C_{k+l}(K)$ in the following fashion:
Let $f=(v_{0},\dots,v_{p-1})$ be a $p$–polygonal face, $(v_{i},v_{j})$ an edge, and $v$ a vertex of $K$, the polygonal cup product is defined for each degree and per each $(k+l)$–cell as:
@@
where $\alpha(i):=\alpha(e_{i})=\alpha(v_{i},v_{(i+1)\%p})$ and %$\%p$ means modulo $p$.
@@
The case $d(\alpha^{0}\cup\beta^{1})$: Let us first simplify the notation and denote $\alpha_{i}=\alpha^{0}(v_{i})$, $\beta_{i}=\beta^{1}(e_{i})$. Let also any index $i$ to be understood as $i\%n$. Let $f=(v_{0},\dots,v_{n-1})$ be an $n$–polygonal face with edges $e_{i}=(v_{i},v_{i+1})$. Due to the indexes being cyclic (modulo $n$), we can reorder the following sums in this manner:
@@
With respect to the announced form of associativity, in the case of the product of three 0–forms we get to multiplying three real numbers on each vertex, which is indeed an associative operation. For the other degrees, we have learned in the proof of Theorem 5.1 that if $K$ is connected and a 0–form $\alpha^{0}$ is constant (which is equivalent to closed), that is $\alpha^{0}(v_{i})=\rho\in\mathbb{R}\;\forall v_{i}\in K$, then the cup product of $\alpha^{0}$ with any $q$–form $\omega^{q}$ is equal to multiplying $\omega^{q}$ by the real number $\rho$. This fact yields:
@@
Concerning the skew–commutativity, looking at the equations (7)–(9) it is easy to see that $\alpha^{0}\cup\beta^{l}=\beta^{l}\cup\alpha^{0}$ for $l=0,1,2$. The case of the product of two 1–forms is slightly more involving. We need to prove that $(\alpha^{1}\cup\beta^{1})(f)=-(\beta^{1}\cup\alpha^{1})(f)$ on an $n$–polygonal face $f=(v_{0},\dots,v_{n-1})$. Let us remember that
@@
where $\alpha(i):=\alpha(e_{i})=\alpha(v_{i},v_{(i+1)\%n})$ and %$\%n$ means modulo $n$. Again, to simplify the notation we will omit writing the symbol %$\%n$ in the indexes and by some index $k$ we will always understand $k\%n$. We next show that

### v4

# A discrete wedge product on general polygonal meshes

@@ 1 Introduction
In this paper, we prove that the wedge product on general polygonal meshes, used in applications in Ptackova and Velho (2021), satisfies the Leibniz product rule with discrete exterior derivative (coboundary operator). Furthermore, we show that it corresponds to a cup product of cochains on orientable 2–dimensional pseudomanifolds, whose highest dimensional cells are any simple polygons. Even though several discretizations of wedge product on simplicial meshes have been studied previously, to the best of our knowledge, similar results for product on general polygonal meshes have not been published previously. We thus partially extend the findings of Schubel et al. (2024) from simplicial to general polygonal setting.
On smooth manifolds, the wedge product allows for building higher degree forms from lower degree ones. Similarly a cup product is a product of two cochains of arbitrary degree $p$ and $q$ that returns a cochain of degree $p+q$ located on (p+q)(p+q)–dimensional cells. The cup product was introduced by J. W. Alexander, E. Čech, and H. Whitney (Whitney, 1957) in 1930’s and it became a well–studied notion in algebraic topology, mainly in the simplicial setting (Munkres, 1984; Fenn, 1983). Later, the cup product was extended also to nn–cubes (Massey, 1991). In (Arnold, 2012) various cup products on simplicial and cubical nn–pseudomanifolds are presented. Arnold (2012) also defines L2 cubical Whitney forms and a cubical cup product that fit together with the wedge product of cubical Whitney forms.
In (Gonzalez-Diaz et al., 2011a) cup products on cubical complexes have been applied to compute cohomology rings of 3D digital images (represented as cubical complexes $Q(I)$), thus simplify the combinatorial structure of $Q(I)$ and obtain a homeomorphic cellular complex $P(I)$ with fewer cells, which improves the computational efficiency. Gonzalez-Diaz et al. (2011a) think of the cup product as the way the nn–dimensional holes obtained in homology are related to each other. It is known that two objects with non isomorphic cohomology rings are not homotopic.
@@ 3 Discrete setting
    Every (n1)(n-1)-cell is a face of at most two $n$-cells.
@@
    such that $e^{n}_{a}=e_{0}^{n},\,e^{n}_{b}=e^{n}_{k}$, and $e^{n}_{i-1}$ and $e^{n}_{i}$ have a common (n1)(n-1)-dimensional face ($i=1,\dots,k$).
If in the condition (2) of above definition, every (n1)(n-1)-dimensional cell is a face of exactly two $n$-cells, then we talk about pseudomanifold without boundary ((Massey, 1991, Chapter IX)). On the other hand, if there exist at least one (n1)(n-1)-cell which is incident with only one $n$-cell, we talk about pseudomanifold with boundary.
@@
Let $K=\{K^{n}\}$ be a pseudomanifold on the topological space $X$. And let $e_{\lambda}^{n}:e_{\mu}^{n-1}]$ be the incidence number of the cells $e_{\lambda}^{n}$ and $e_{\mu}^{n-1}$, for $n>0$, such that:
@@
    If $e_{\mu}^{n-1}$ is not a face of $e_{\lambda}^{n}$, then $e_{\lambda}^{n}:e_{\mu}^{n-1}]=0$.
@@
    If $e_{\mu}^{n-1}$ is a face of $e_{\lambda}^{n}$, then $e_{\lambda}^{n}:e_{\mu}^{n-1}]=\pm 1$.
@@
    If $e_{\alpha}^{0}$ and $e_{\beta}^{0}$ are the two vertices of the 1-cell $e_{\lambda}^{1}$, then $e_{\lambda}^{1}:e_{\alpha}^{0}]+[e_{\lambda}^{1}:e_{\beta}^{0}]=0$.
@@
    Let $e_{\lambda}^{n}$ and $e_{\rho}^{n-2}$ be cells such that $e_{\rho}^{n-2}\prec e_{\lambda}^{n}$; let $e_{\alpha}^{n-1}$ and $e_{\beta}^{n-1}$ denote the unique (n1)(n-1)-cells $e^{n-1}$ such that $e_{\rho}^{n-2}\prec e^{n-1}\prec e_{\lambda}^{n-1}$. Then
    $e_{\lambda}^{n}:e_{\alpha}^{n-1}][e_{\alpha}^{n-1}:e_{\rho}^{n-2}]+[e_{\lambda}^{n}:e_{\beta}^{n-1}][e_{\beta}^{n-1}:e_{\rho}^{n-2}]=0.$
@@
Let $K$ be an $n$-dimensional pseudomanifold (possibly with or without boundary), and let $e_{1}^{n},e_{2}^{n}$ be $n$-cells with a common (n1)(n-1)-cell $e^{n-1}$. We define orientations for $e_{1}^{n}$ and $e_{2}^{n}$ to be coherent (with respect to $e^{n-1}$) if:
$e_{1}^{n}:e^{n-1}]+[e_{2}^{n}:e^{n-1}]=0.$
A set of orientations for all the $n$-cells of $K$ is said to be coherent if it is coherent in the above sense for any pair of $n$-cells with a common (n1)(n-1)-face.
$K$ is said to be orientable if all its $n$-cells can be oriented such that any pair of $n$-cells sharing an (n1)(n-1)-dimensional face are oriented coherently. Otherwise it is called nonorientable.
@@
$0\longrightarrow C_{n}(K)\overset{\partial_{n}}{\longrightarrow}\cdots\overset{\partial_{k+1}}{\longrightarrow}C_{k}(K)\overset{\partial_{k}}{\longrightarrow}\cdots\overset{\partial_{1}}{\longrightarrow}C_{0}(K)\longrightarrow 0.$
@@ 4 Discrete Differential Forms and the Exterior Derivative
Figure 1: Boundary homomorphism and exterior derivative on a polygonal complex. In this figure, our pseudomanifold $K$ has 2–chain $C_{2}=(f_{0},f_{1})$, 1–chain $C_{1}=(e_{0},\dots,e_{5}$), and 0–chain $C_{0}=(v_{0},\dots,v_{4})$. If 2–form $\gamma^{2}$ was given as the exterior derivative of 1–form $\beta^{1}$, then $\gamma=d\beta$ would be defined per faces $f_{0},f_{1}$ by equation (2). Concretely $\gamma(f_{1})=d\beta(f_{1})$ is:
@@
$0\overset{d_{n}}{\longleftarrow}C^{n}(K)\overset{d_{n-1}}{\longleftarrow}\cdots\overset{d_{k}}{\longleftarrow}C^{k}(K)\overset{d_{k-1}}{\longleftarrow}\cdots\overset{d_{0}}{\longleftarrow}C^{0}(K)\longleftarrow 0.$
A qq–form $\alpha$ is said to be a closed form (cocycle) if $d\alpha=0$, and it is called an exact form (coboundary) if there exist a (q1)(q-1)–form $\beta$ such that $d\beta=\alpha$.
Note that each exact $q$-form $\alpha=d\beta$ is closed, since $d\alpha=dd\beta=0$. Two closed $q$-forms are cohomologous if they differ by an exact qq–form. We denote cohomology classes by $H^{\ast}$:
@@ 5 Wedge and Cup Products
A differential $k$-form $\alpha$ on a smooth nn–manifold $M$ is a tensor field of type $(0,k)$ that is completely antisymmetric, i.e., for $\alpha\in\Omega^{k}(M)\subset T^{0}_{k}(M)$ we have $\alpha(v,w)=-\alpha(w,v)$ for all $v,w\in TM$. Let $\{e_{1},\dots,e_{n}\}$ be a basis of the tangent space $T_{x}M$ to $M$ at point $x\in M$. Let $\alpha\in T^{0}_{k}(M)$ and $\beta\in T^{0}_{l}(M)$, we define their wedge product $\alpha\wedge\beta\in T^{0}_{k+l}(M)$ at the point $x\in M$ by
@@ 5.1 Simplicial and cubical cup products
Let us first recall the definition of simplicial Whitney forms. Let $c=(v_{0},\dots,v_{p})$ be a pp–simplex with barycentric coordinates $x_{0},\dots,x_{p}$. Thus $0\leq x_{j}\leq 1\;\forall j$ and $\sum_{i=0}^{p}=1$. The simplicial Whitney pp–form corresponding to the simplex $c$ is
@@
where $c$ is an (p+q)(p+q)–simplex and $W$ is the simplicial Whitney form.
@@
Let $(v_{0},\dots,v_{k})$ be a kk–simplex of a pseudomanifold $K$, the simplicial cup product $\cup:C^{p}(K,\mathbb{R})\times C^{q}(K,\mathbb{R})\rightarrow C^{p+q}(K,\mathbb{R})$, i.e., the cup product defined over each simplex $(v_{0},\dots,v_{k})$ reads:
@@
In (Arnold, 2012, Definition 3.2.1) the author defines a cubical cup product $\cup_{c}$, for which the product of two cochains agrees with the wedge product of their cubical Whitney forms (see (Arnold, 2012, Section 3.2.2)), thus the Whitney map provides a connection of cubical cohomology with de Rham cohomology ((Arnold, 2012, Theorem 3.2.12)). From (Arnold, 2012, Definition 3.2.1) we computed the cup product on kk–dimensional cubes for $0\leq k\leq 2$ and we present the formulas in the following definition:
@@
The cubical cup product $\cup:C^{p}(K,\mathbb{R})\times C^{q}(K,\mathbb{R})\rightarrow C^{p+q}(K,\mathbb{R})$ on a quadrilateral pseudomanifold $K$ is defined by its action on a kk–dimensional cube $(v_{0},\dots,v_{2^{k}-1})$ as:
@@ 5.2 The cup product on polygons
As advertised earlier, we can extend the definitions of the discrete wedge product to any pp–polygon such that it still satisfies the defining properties of a cup product (Definition 5.1) and therefore automatically also the properties of Proposition 5.2. Moreover, we show that our polygonal cup product is a skew–commutative bilinear operation on any polygonal forms, not just on closed ones, viz. the Proposition 5.3.
Let us remember that a discrete qq–form $\alpha^{q}$ is a qq–cochain defined per each element of a qq–chain $C_{q}(K)$:
@@
Furthermore, we sum two discrete qq–forms $\alpha^{q},\beta^{q}$ by summing their values per each element of a qq–chain $C_{q}(K)$ and multiply a qq–form $\alpha^{q}$ by a scalar $t\in\mathbb{R}$ in element–wise fashion too:
@@
Let $K$ be a 2–pseudomanifold whose 2–cells (faces) are polygons. The cup product $C^{k}(K,\mathbb{R})\times C^{l}(K,\mathbb{R})\rightarrow C^{k+l}(K,\mathbb{R})$ of two discrete forms $\alpha^{k},\beta^{l}$ is a (k+l)(k+l)–form $\alpha^{k}\cup\beta^{l}$ defined on each (k+l)(k+l)–cell of $C_{k+l}(K)$ in the following fashion:
Let $f=(v_{0},\dots,v_{p-1})$ be a pp–polygonal face, $(v_{i},v_{j})$ an edge, and $v$ a vertex of $K$, the polygonal cup product is defined for each degree and per each (k+l)(k+l)–cell as:
@@
where $\alpha(i):=\alpha(e_{i})=\alpha(v_{i},v_{(i+1)\%p})$ and $\%p$ means modulo $p$.
@@
The case $d(\alpha^{0}\cup\beta^{1})$: Let us first simplify the notation and denote $\alpha_{i}=\alpha^{0}(v_{i})$, $\beta_{i}=\beta^{1}(e_{i})$. Let also any index $i$ to be understood as $i\%n$. Let $f=(v_{0},\dots,v_{n-1})$ be an nn–polygonal face with edges $e_{i}=(v_{i},v_{i+1})$. Due to the indexes being cyclic (modulo $n$), we can reorder the following sums in this manner:
@@
With respect to the announced form of associativity, in the case of the product of three 0–forms we get to multiplying three real numbers on each vertex, which is indeed an associative operation. For the other degrees, we have learned in the proof of Theorem 5.1 that if $K$ is connected and a 0–form $\alpha^{0}$ is constant (which is equivalent to closed), that is $\alpha^{0}(v_{i})=\rho\in\mathbb{R}\;\forall v_{i}\in K$, then the cup product of $\alpha^{0}$ with any qq–form $\omega^{q}$ is equal to multiplying $\omega^{q}$ by the real number $\rho$. This fact yields:
@@
Concerning the skew–commutativity, looking at the equations (7)–(9) it is easy to see that $\alpha^{0}\cup\beta^{l}=\beta^{l}\cup\alpha^{0}$ for $l=0,1,2$. The case of the product of two 1–forms is slightly more involving. We need to prove that $(\alpha^{1}\cup\beta^{1})(f)=-(\beta^{1}\cup\alpha^{1})(f)$ on an nn–polygonal face $f=(v_{0},\dots,v_{n-1})$. Let us remember that
@@
where $\alpha(i):=\alpha(e_{i})=\alpha(v_{i},v_{(i+1)\%n})$ and $\%n$ means modulo $n$. Again, to simplify the notation we will omit writing the symbol $\%n$ in the indexes and by some index $k$ we will always understand $k\%n$. We next show that